# Imaging Genetics Analysis of Alzheimer’s Disease

### v1

@@ Appendix A Appendix section
Characteristic Cognitively Normal (417) Early Mild CI (310) Late Mild CI (562) Diagnosed AD (342) Demographics Sex of Patient, n (%)    Male 209 (50.12%) 171 (55.16%) 344 (61.21%) 189 (55.26%)    Female 208 (49.88%) 139 (44.84%) 218 (38.79%) 153 (44.74%) Race of Patient, n (%)    White 376 (90.17%) 286 (92.26%) 526 (93.59%) 317 (92.69%)    Black 30 (7.19%) 8 (2.58%) 22 (3.91%) 14 (4.09%)    Other 11 (2.64%) 16 (5.16%) 14 (2.42%) 11 (3.22%) Formal Years of Education, Mean (SD) 16.28 (2.73) 15.96 (2.66) 15.88 (2.94) 15.18 (2.99) Risk Factors Baseline Age (years), Mean (SD) 74.76 (5.73) 71.19 (7.50) 73.99 (7.50) 75.03 (7.79) (Apolipoprotein E4) Genetic Status, n (%)    Absent 301 (72.53%) 175 (57.19%) 256 (45.71%) 113 (33.43%)    Present 114 (27.47%) 131 (42.81%) 304 (54.29%) 225 (66.57%) Cognitive Exams ADAS (13), Mean (SD) 9.34 (4.32) 12.65 (5.42) 18.66 (6.52) 29.87 (8.05) CDRSB, Mean (SD) 0.03 (0.13) 1.29 (0.76) 1.65 (0.92) 4.39 (1.67) MMSE, Mean (SD) 29.07 (1.12) 28.34 (1.56) 27.18 (1.80) 23.22 (2.07) Rey Auditory General Test, Mean (SD) 44.35 (9.84) 39.55 (10.71) 31.32 (9.51) 22.82 (7.55) Cerebrospinal Fluid (CSF) Measurements Amyloid-Beta levels, Mean (SD) 1,327.68 (660.95) 1,178.19 (587.49) 889.14 (490.26) 691.44 (416.43) Tau levels, Mean (SD) 238.45 (88.96) 256.41 (121.74) 308.95 (128.96) 367.84 (144.92) Phosphorylated Tau levels, Mean (SD) 22.00 (9.08) 24.25 (13.69) 30.46 (14.66) 36.66 (15.76)

### v2

@@ Appendix A Appendix section
Characteristic Cognitively Normal (417) Early Mild CI (310) Late Mild CI (562) Diagnosed AD (342)
Demographics
Sex of Patient, n (%)
     Male 209 (50.12%) 171 (55.16%) 344 (61.21%) 189 (55.26%)
     Female 208 (49.88%) 139 (44.84%) 218 (38.79%) 153 (44.74%)
Race of Patient, n (%)
     White 376 (90.17%) 286 (92.26%) 526 (93.59%) 317 (92.69%)
     Black 30 (7.19%) 8 (2.58%) 22 (3.91%) 14 (4.09%)
     Other 11 (2.64%) 16 (5.16%) 14 (2.42%) 11 (3.22%)
Formal Years of Education, Mean (SD) 16.28 (2.73) 15.96 (2.66) 15.88 (2.94) 15.18 (2.99)
Risk Factors
Baseline Age (years), Mean (SD) 74.76 (5.73) 71.19 (7.50) 73.99 (7.50) 75.03 (7.79)
(Apolipoprotein E4) Genetic Status, n (%)
     Absent 301 (72.53%) 175 (57.19%) 256 (45.71%) 113 (33.43%)
     Present 114 (27.47%) 131 (42.81%) 304 (54.29%) 225 (66.57%)
Cognitive Exams
ADAS (13), Mean (SD) 9.34 (4.32) 12.65 (5.42) 18.66 (6.52) 29.87 (8.05)
CDRSB, Mean (SD) 0.03 (0.13) 1.29 (0.76) 1.65 (0.92) 4.39 (1.67)
MMSE, Mean (SD) 29.07 (1.12) 28.34 (1.56) 27.18 (1.80) 23.22 (2.07)
Rey Auditory General Test, Mean (SD) 44.35 (9.84) 39.55 (10.71) 31.32 (9.51) 22.82 (7.55)
Cerebrospinal Fluid (CSF) Measurements
Amyloid-Beta levels, Mean (SD) 1,327.68 (660.95) 1,178.19 (587.49) 889.14 (490.26) 691.44 (416.43)
Tau levels, Mean (SD) 238.45 (88.96) 256.41 (121.74) 308.95 (128.96) 367.84 (144.92)
Phosphorylated Tau levels, Mean (SD) 22.00 (9.08) 24.25 (13.69) 30.46 (14.66) 36.66 (15.76)